# AuON: A Linear-time Alternative to Orthogonal Momentum Updates

### v3

@@ Abstract
Orthogonal momentum gradient updates have emerged to overcome the limitations of vector-based optimizers like Adam. The vector-based optimizer Adam suffers from high memory costs and ill-conditioned momentum gradient updates. However, traditional Orthogonal momentum approaches, such as SVD/QR decomposition, suffer from high computational and memory costs and underperform compared to well-tuned SGD with momentum.Recent advances, such as Muon, improve efficiency by applying momentum before orthogonalization and approximate orthogonal matrices via Newton-Schulz iterations, which gives better GPU utilization, active high TFLOPS, and reduces memory usage by up to 3x. Nevertheless, Muon(Vanilla) suffers from exploding attention logits and has cubic computation complexity. In this paper we , deep dive into orthogonal momentum gradient updates to find the main properties that help Muon to achieve remarkable performance.We propose AuON (Alternative Unit-norm momentum updates by Normalized nonlinear scaling), a linear-time optimizer that achieves strong performance without approximate orthogonal matrices, while preserving structural alignment and reconditioning ill-posed updates. AuON has an automatic (”emergency brake”) to handle exploding attention logits.. We further introduce a hybrid variant ( Hybrid-AuON)that applies the linear transformations with Newton-Schulz iterations which out performs Muon in the language modeling tasks. Code is available at: https://github.com/ryyzn9/AuON
@@ 1 Introduction
The pursuit of computational efficiency in training Large Language Models (LLMs) has catalyzed a shift away from purely scalar, coordinate-wise adaptive optimizers toward geometry-aware methods. For over a decade, the optimization landscape has been hegemonized by first-order adaptive moment estimation algorithms, principally Adam and its decoupled weight-decay variant, AdamW. These methods, while robust, operate under the simplifying assumption that the parameter space is Euclidean and that each parameter can be adapted independently based on its gradient history. However, as neural architectures have scaled into the regime of trillions of parameters—exemplified by models like the Mixture-of-Experts (MoE) limitations of this coordinate-wise paradigm have become increasingly stark. The inefficiencies manifest not merely in convergence speed but in the disconnect between the optimizer’s update steps and the underlying matrix structure of modern neural networks. Principally, Adam and its decoupled weight-decay variant, AdamW, suffer from ill-conditioning of gradient and momentum updates due to the momentum for a linear layer, naturally, a 2D matrix tends to become almost low-rank in practice.This means that only a small number of dominant directions really drive the updates. Though rare directions seem minor, they are often essential for effective learning and can help capture more nuanced patterns in the data. Empirically, these momentum updates often exhibit a high condition number, with most of the energy concentrated in a few dominant directions. In practical terms, the update vectors are nearly low-rank: a handful of directions dictate the optimization trajectory while many potentially informative directions may be suppressed. This imbalance reminds us of a squashed ball that can only roll efficiently along a single axis, ignoring other pathways that may be equally important for generalization and representation learning. One solution is to make all update directions unit length; recent work has proposed orthogonalization of gradients and momentum updates to achieve this property. Adam and its variant AdamW have another critical limitation: they require keeping two extra variables for every model parameter, meaning the optimizer state takes up about twice as much memory as the model itself. Furthermore, Adam treats all parameters as a single long vector, updating each value independently without considering any internal structure of the parameters.
@@
tuddenham2022orthogonalisinggradientsspeedneural proposed an approach for neural network optimization in which the gradient is first orthogonalized via singular value decomposition (SVD), followed by the application of momentum, and then the resulting momentum term is used as the update. They refer to this method as Orthogonal-SGDM. In their experiments, they observed that even in the best-performing configuration, Orthogonal-SGDM was outperformed by a well-tuned standard SGD with momentum. This is because applying momentum after orthogonalization which damages the momentum mechanism: orthogonalizing gradients before accumulation prevents momentum from effectively reducing variance and maintaining beneficial directional information. Moreover, orthogonality erases singular-value magnitudes and over-constrains the step, collapsing its singular-value structure to an isometry. In effect, the update becomes a spectral-norm-constrained move that discards useful magnitude information, wiping out correlations between update directions. Making all updates to unit length may also increase harmful alignment effects. Recent advances, such as Muon jordan2024muon, improve efficiency and performance by approximating an orthogonal matrix using Newton–Schulz iterations rather than a full orthogonal matrix using SVD, and by reordering the momentum update to occur before orthogonalization. This makes the Muon computational efficient in GPU, reduces the wall-clock time, and gives much higher throughput, though Muon has cubic complexity, is 2 x faster than Adam and AdamW, but still suffers from exploding attention logits. We find out that, by bounding the momentum updates under a spectral-norm trust region can preserve the directional information without requiring explicit orthogonalization
In this paper, we focus on developing an alternative approach to bound updates with high condition numbers under a unit-norm constraint, without explicit Orthogonalization. Our goal is to achieve strong performance with without compromising efficiency or speed. Empirically, we find that normalization followed by a hyperbolic-cosine scaling transformation can bound the updates with high condition numbers under spectral-norm trust region and uses an automatic ”emergency brake” to significantly more resistant to exploding attention logits than Muon, especially in heavy‑tailed regimes.
@@ 3 Methods
We hypothesize that forcing all update directions to unit length can be problematic, as not all directions contribute equally to optimization progress some may be harmful (having a negative impact) or irrelevant to loss reduction.At first glance, it seems impossible: how can one suppress ”harmful” directions without knowing what they are in advance? If one never explicitly decorrelates or identifies bad modes , how does it preferentially suppress them?
@@
Our goal is to develop a method that suppress the harmful directions(spiky-gradient updates) or alignments and preserves the beneficial properties of orthogonalization while selectively scaling directions under unit-norm: one solution is to decrease the scales of rare update directions relative to dominant ones and keep them all under a unit-norm trust region, meaning prioritizing directions with favorable conditioning that correspond to well-conditioned subspaces of the loss landscape under a spectral-norm trust-region.To do that we can apply a temperature-scaled softmax update matrix, followed by L2-renormalization, to bound the step under a trust-region. But computing softmax may be problematic as it introduces computational bottlenecks and it can also dampens useful directions as a side effect of dampening which is crucial for effective learning. We empirically find that the normalization with hyperbolic functions (cosh ) helps us bound the momentum updates under spectral-norm trust-region kovalev2025understandinggradientorthogonalizationdeep and it is a clever statistical mechanism (the cosh-RMS scaling) that automatically dampens harmful directions as a side effect of dampening all spiky components as a result we get ”emergency brake” property which can handle exploding attention logits than Muon, especially in heavy-tailed regimes.”Emergency Brake” is an emergent property of the hyperbolic cosine (cosh) normalization step.
@@ 3.1 Nonlinear reshaping via hyperbolic cosine RMS scaling
Our main goal is to keep all the updated directions under unit spectral norm and remove the harmful directions. We empirically find out that the updated matrix divided by the scale factor of the RMS magnitude of cosh() helps us bound the dominant update directions that have a high condition number under unit spectral norm, and helps us to preserve the orthogonal-like properties. By doing this, we remove the harmful alignment and stabilize the updates and equitable learning dynamics. The overall equation is
@@
This encourages a spread of activations (diversity) without enforcing strict orthogo- nality Peletier_2023 and X is the updated momentum vector direction
@@
Define $\mathrm{rms}:=\|\cosh(\mathrm{update})\|_{F}/\sqrt{N}$, where $\cosh$ is applied only to compute a global, tail-sensitive scale article. Because $\cosh$ is even and rapidly increasing in $|x|$, heavy tails inflate $\mathrm{rms}$, which reduces the overall step size when forming $U:=\mathrm{update}/(\mathrm{rms}+10^{-8})$. Crucially, $\cosh$ is not applied to the propagated vector: $U$ is a uniform rescaling of $\mathrm{update}$, so the signs and all relative component ratios of $\mathrm{update}$ are preserved in $U$. This yields scale invariance with tail-aware damping, without introducing per-coordinate reweighting in the final update. As a resust as soon as attention logits start to explode (creating gradient spikes), AuON automatically suppresses the update to near-zero. It effectively freezes the unstable layer rather than pushing it further.it is tail-aware. It uses the cosh function to detect when the optimization landscape is becoming unstable (spiky gradients) and reacts by shrinking the step size
@@ 3.2 Hybrid Approach
Discussion. Hybrid-AuON is not a full semi-orthogonalization method. It preserves part of the correlation structure but guarantees bounded, tail-aware updates. Its purpose is to offer a middle ground: more stable than plain AuON, but cheaper than Muon.
@@ 3.3 Temperature–Scaled AuON: Sensitivity to Tensor Size and Heavy Tails
For small arguments, $\cosh(x)\approx 1+x^{2}/2$, giving $s\approx 1+\mathbb{E}[g_{\text{scaled}}^{2}],$ but since $g_{\text{scaled}}$ has variance proportional to $N^{2\alpha}$, even the quadratic term grows with tensor size. For rare large entries, $\cosh(x)\sim\tfrac{1}{2}e^{|x|}$, so a tiny number of outliers with $|x|\gg 1$ dominate $s$ exponentially. Raising $s$ to the exponent $3.5/2$ further super-linearizes this effect: if tail magnitude doubles, the brake factor $r$ grows by more than a factor of two, causing the step norm to collapse sharply.
@@ 3.4 Spectral Trust-Region Guarantee and Reduction of Correlation Energy in AuON
Hence large spikes are suppressed at an exponential rate:
@@ 3.5 Emergent “Emergency Brake” Behavior in AuON via Hyperbolic-Cosine Normalization
A critical failure mode in large-scale Transformer training is the phenomenon of “exploding at tention logits.” As weight norms growunbounded, thequery-keydotproductsdiverge, causing Softmax saturation. This results in gradient matrices that are geometrically sparse and heavy tailed (high kurtosis).
AuON introduces a hyperbolic-cosine–based normalization that acts as an automatic “emergency brake” when gradients become unstable, for example under exploding attention heads. Unlike Muon, which normalizes to unit norm regardless of the gradient distribution, AuON computes a scalar factor
@@ 3.5.4 Empirical Verification at Production Scale (d=4096)
To stress-test the “Emergency Brake” mechanism at a realistic scale, we constructed a Transformer layer with model dimension d = 4096 and an MLP expansion factor of 4(Dmlp = 4096), totaling approximately 100.6 Million million parameters. Crucially, all normalization layers (LayerNorm/RM SNorm) were removed and the learning rate was kept at 0.24 . This architecture creates a high-pressure optimization environment where weight accumulation leads directly to attention logit explosion. We compared the Newton-Schulz semi-orthogonal update (Muon) against AuON (with temperature scaling )over 1000 steps.
Figure 3: The Emergency Brake in Action. The Muon optimizer (Red) exhibits step-function explosions, driving logits to hard saturation (greater than 30). In contrast, AuON (Blue) and AuON+5(Ns=5)(Green) demonstrat the theoretically predicted logarithmic braking curve, capping logit growth at approximately 12 even after 1,000 steps of unnormalized accumulation.
@@
  • AuON and Hybrid-AuON(Ns=5) (Brake Engaged): The blue and green trajectorys confirms the exponential suppression derived in Eq. (9). As logits attempt to grow, the gradient kurtosis rises, triggering the $\cosh$ term to inflate the denominator $r$. The result is a smooth, controlled saturation curve rather than a catastrophic explosion.
@@ 3.5.5 Long-Horizon Stability: The Dissipative Property
Figure 4: Thermodynamics of Optimization (20k Steps). The Muon optimizer (Red) exhibits high-variance chaotic behavior with a persistent upward drift, indicative of unregulated energy injection. In stark contrast, AuON (Blue) demonstrates a dissipative trajectory. After an initial rise due to the lack of normalization, the self-regulating cosh mechanism forces a inflection point at step approximately 2,500, after which the logit magnitudes steadily decay. This proves that AuON actively removes instability energy from the system over time.
@@ 4.2 Comparison of Hybrid-AuON and Muon on FineWebEDU-10B
We conduct a controlled paired comparison between Temperature-scaled Hybrid-AuON and Muon on a H100 gpu with a 12-layer GPT-style decoder with hidden size $d=768$ and 6 attention heads(168 Million parameter). Training data are streamed from the 10B-token FineWebEDU kydlicek2025finepdfs corpus using a custom generator that samples contiguous 8k-token sequences while respecting document boundaries via alignment to the GPT-2 end-of-sequence token (50256). Each training step consumes a single 8k-token sequence for both input and target. Validation uses the same sequence length with a fixed budget of 8 sequences (64k tokens), evaluated every 100 steps.
@@
yielding super-linear suppression that engages earlier under spiky gradients.It just change how heard and when the brack engages

### v4

@@ Abstract
Orthogonal momentum gradient updates have emerged to overcome the limitations of vector-based optimizers like Adam. The vector-based optimizer Adam suffers from high memory costs and ill-conditioned momentum gradient updates. However, traditional Orthogonal momentum approaches, such as SVD/QR decomposition, suffer from high computational and memory costs and underperform compared to well-tuned SGD with momentum. Recent advances, such as Muon, improve efficiency by applying momentum before orthogonalization and approximate orthogonal matrices via Newton-Schulz iterations, which gives better GPU utilization, active high TFLOPS, and reduces memory usage by up to 3x. Nevertheless, Muon(Vanilla) suffers from exploding attention logits and has cubic computation complexity. In this paper we deep dive into orthogonal momentum gradient updates to find the main properties that help Muon achieve remarkable performance. We propose AuON (Alternative Unit-norm momentum updates by Normalized nonlinear scaling), a linear-time optimizer that achieves strong performance without approximate orthogonal matrices, while preserving structural alignment and reconditioning ill-posed updates. AuON has an automatic (”emergency brake”) to handle exploding attention logits. We further introduce a hybrid variant ( Hybrid-AuON)that applies the linear transformations with Newton-Schulz iterations, which outperforms Muon in the language modeling tasks. Code is available at: https://github.com/ryyzn9/AuON
@@ 1 Introduction
The pursuit of computational efficiency in training Large Language Models (LLMs) has catalyzed a shift away from purely scalar, coordinate-wise adaptive optimizers toward geometry-aware methods. For over a decade, the optimization landscape has been hegemonized by first-order adaptive moment estimation algorithms, principally Adam and its decoupled weight-decay variant, AdamW. These methods, while robust, operate under the simplifying assumption that the parameter space is Euclidean and that each parameter can be adapted independently based on its gradient history. However, as neural architectures have scaled into the regime of trillions of parameters—exemplified by models like the Mixture-of-Experts (MoE) limitations of this coordinate-wise paradigm have become increasingly stark. The inefficiencies manifest not merely in convergence speed but in the disconnect between the optimizer’s update steps and the underlying matrix structure of modern neural networks. Principally, Adam and its decoupled weight-decay variant, AdamW, suffer from ill-conditioning of gradient and momentum updates due to the momentum for a linear layer; naturally, a 2D matrix tends to become almost low-rank in practice. This means that only a small number of dominant directions really drive the updates. Although rare directions seem minor, they are often essential for effective learning and can help capture more nuanced patterns in the data. Empirically, these momentum updates often exhibit a high condition number, with most of the energy concentrated in a few dominant directions. In practical terms, the update vectors are nearly low-rank: a handful of directions dictate the optimization trajectory while many potentially informative directions may be suppressed. This imbalance reminds us of a squashed ball that can only roll efficiently along a single axis, ignoring other pathways that may be equally important for generalization and representation learning. One solution is to make all update directions unit length; recent work has proposed orthogonalization of gradients and momentum updates to achieve this property. Adam and its variant AdamW have another critical limitation: they require keeping two extra variables for every model parameter, meaning the optimizer state takes up about twice as much memory as the model itself. Furthermore, Adam treats all parameters as a single long vector, updating each value independently without considering any internal structure of the parameters.
@@
tuddenham2022orthogonalisinggradientsspeedneural proposed an approach for neural network optimization in which the gradient is first orthogonalized via singular value decomposition (SVD), followed by the application of momentum, and then the resulting momentum term is used as the update. They refer to this method as Orthogonal-SGDM. In their experiments, they observed that even in the best-performing configuration, Orthogonal-SGDM was outperformed by a well-tuned standard SGD with momentum. This is because applying momentum after orthogonalization damages the momentum mechanism: orthogonalizing gradients before accumulation prevents momentum from effectively reducing variance and maintaining beneficial directional information. Moreover, orthogonality erases singular-value magnitudes and over-constrains the step, collapsing its singular-value structure to an isometry. In effect, the update becomes a spectral-norm-constrained move that discards useful magnitude information, wiping out correlations between update directions. Making all updates to unit length may also increase harmful alignment effects. Recent advances, such as Muon jordan2024muon, improve efficiency and performance by approximating an orthogonal matrix using Newton–Schulz iterations rather than a full orthogonal matrix using SVD, and by reordering the momentum update to occur before orthogonalization. This makes the Muon computational efficient in GPU, reduces the wall-clock time, and gives much higher throughput, though Muon has cubic complexity, is 2 x faster than Adam and AdamW, but still suffers from exploding attention logits. We find out that, by bounding the momentum updates under a spectral-norm trust region can preserve the directional information without requiring explicit orthogonalization
In this paper, we focus on developing an alternative approach to bound updates with high condition numbers under a unit-norm constraint, without explicit Orthogonalization. Our goal is to achieve strong performance with without compromising efficiency or speed. Empirically, we find that normalization followed by a hyperbolic-cosine scaling transformation can bound the updates with high condition numbers under the spectral-norm trust region and uses an automatic ”emergency brake” to be significantly more resistant to exploding attention logits than Muon, especially in heavy‑tailed regimes.
@@ 3 Methods
We hypothesize that forcing all update directions to unit length can be problematic, as not all directions contribute equally to optimization progress, some may be harmful (having a negative impact), or irrelevant to loss reduction.At first glance, it seems impossible: how can one suppress ”harmful” directions without knowing what they are in advance? If one never explicitly decorrelates or identifies bad modes, how does it preferentially suppress them?
@@
Our goal is to develop a method that suppress the harmful directions(spiky-gradient updates) or alignments and preserves the beneficial properties of orthogonalization while selectively scaling directions under unit-norm: one solution is to decrease the scales of rare update directions relative to dominant ones and keep them all under a unit-norm trust region, meaning prioritizing directions with favorable conditioning that correspond to well-conditioned subspaces of the loss landscape under a spectral-norm trust-region.To do that we can apply a temperature-scaled softmax update matrix, followed by L2-renormalization, to bound the step under a trust-region. But computing softmax may be problematic as it introduces computational bottlenecks and it can also dampens useful directions as a side effect of dampening, which is crucial for effective learning. We empirically find that the normalization with hyperbolic functions (cosh ) helps us bound the momentum updates under spectral-norm trust-region kovalev2025understandinggradientorthogonalizationdeep and it is a clever statistical mechanism (the cosh-RMS scaling) that automatically dampens harmful directions as a side effect of dampening all spiky components. As a result we get ”emergency brake” property which can handle exploding attention logits than Muon, especially in heavy-tailed regimes.”Emergency Brake” is an emergent property of the hyperbolic cosine (cosh) normalization step.
@@ 3.1 Nonlinear reshaping via hyperbolic cosine RMS scaling
Our main goal is to keep all the updated directions under the unit spectral norm and remove the harmful directions. We empirically find out that the updated matrix divided by the scale factor of the RMS magnitude of cosh() helps us bound the dominant update directions that have a high condition number under unit spectral norm, and helps us to preserve the orthogonal-like properties. By doing this, we remove the harmful alignment and stabilize the updates and equitable learning dynamics. The overall equation is
@@
This encourages a spread of activations (diversity) without enforcing strict orthogo- nality Peletier_2023, and X is the updated momentum vector direction
@@
Define $\mathrm{rms}:=\|\cosh(\mathrm{update})\|_{F}/\sqrt{N}$, where $\cosh$ is applied only to compute a global, tail-sensitive scale article. Because $\cosh$ is even and rapidly increasing in $|x|$, heavy tails inflate $\mathrm{rms}$, which reduces the overall step size when forming $U:=\mathrm{update}/(\mathrm{rms}+10^{-8})$. Crucially, $\cosh$ is not applied to the propagated vector: $U$ is a uniform rescaling of $\mathrm{update}$, so the signs and all relative component ratios of $\mathrm{update}$ are preserved in $U$. This yields scale invariance with tail-aware damping, without introducing per-coordinate reweighting in the final update. As a result, as soon as attention logits start to explode (creating gradient spikes), AuON automatically suppresses the update to near-zero. It effectively freezes the unstable layer rather than pushing it further.it is tail-aware. It uses the cosh function to detect when the optimization landscape is becoming unstable (spiky gradients) and reacts by shrinking the step size
@@ 3.2 Hybrid Approach
Discussion. Hybrid-AuON is not a full semi-orthogonalization method. It preserves part of the correlation structure, but guarantees bounded, tail-aware updates. Its purpose is to offer a middle ground: more stable than plain AuON, but cheaper than Muon.
@@ 3.3 Temperature–Scaled AuON: Sensitivity to Tensor Size and Heavy Tails
For small arguments, $\cosh(x)\approx 1+x^{2}/2$, giving $s\approx 1+\mathbb{E}[g_{\text{scaled}}^{2}],$ but since $g_{\text{scaled}}$ has variance proportional to $N^{2\alpha}$, even The quadratic term grows with tensor size. For rare large entries, $\cosh(x)\sim\tfrac{1}{2}e^{|x|}$, so a tiny number of outliers with $|x|\gg 1$ dominate $s$ exponentially. Raising $s$ to the exponent $3.5/2$ further super-linearizes this effect: If the tail magnitude doubles, the brake factor $r$ grows by more than a factor of two, causing the step norm to collapse sharply.
@@ 3.4 Spectral Trust-Region Guarantee and Reduction of Correlation Energy in AuON
Hence, large spikes are suppressed at an exponential rate:
@@ 3.5 Emergent “Emergency Brake” Behavior in AuON via Hyperbolic-Cosine Normalization
A critical failure mode in large-scale Transformer training is the phenomenon of “exploding at tention logits.” As weight norms grow unbounded, the query-key dot products diverge, causing Softmax saturation. This results in gradient matrices that are geometrically sparse and heavy-tailed (high kurtosis).
AuON introduces a hyperbolic-cosine-based–based normalization that acts as an automatic “emergency brake” when gradients become unstable, for example under exploding attention heads. Unlike Muon, which normalizes to unit norm regardless of the gradient distribution, AuON computes a scalar factor
@@ 3.5.4 Empirical Verification at Production Scale (d=4096)
To stress-test the “Emergency Brake” mechanism at a realistic scale, we constructed a Transformer layer with model dimension d = 4096 and an MLP expansion factor of 4(Dmlp = 4096), totaling approximately 100.6 million parameters. Crucially, all normalization layers (LayerNorm/RM SNorm) were removed and the learning rate was kept at 0.24 . This architecture creates a high-pressure optimization environment where weight accumulation leads directly to attention logit explosion. We compared the Newton-Schulz semi-orthogonal update (Muon) against AuON (with temperature scaling )over 1000 steps.
Figure 3: The Emergency Brake in Action. The Muon optimizer (Red) exhibits step-function explosions, driving logits to hard saturation (greater than 30). In contrast, AuON (Blue) and AuON+5(Ns=5)(Green) demonstrate the theoretically predicted logarithmic braking curve, capping logit growth at approximately 12 even after 1,000 steps of unnormalized accumulation.
@@
  • AuON and Hybrid-AuON(Ns=5) (Brake Engaged): The blue and green trajectories confirm the exponential suppression derived in Eq. (9). As logits attempt to grow, the gradient kurtosis rises, triggering the $\cosh$ term to inflate the denominator $r$. The result is a smooth, controlled saturation curve rather than a catastrophic explosion.
@@ 3.5.5 Long-Horizon Stability: The Dissipative Property
Figure 4: Thermodynamics of Optimization (20k Steps). The Muon optimizer (Red) exhibits high-variance chaotic behavior with a persistent upward drift, indicative of unregulated energy injection. In stark contrast, AuON (Blue) demonstrates a dissipative trajectory. After an initial rise due to the lack of normalization, the self-regulating cosh mechanism forces an inflection point at approximately step 2,500, after which the logit magnitudes steadily decay. This proves that AuON actively removes instability energy from the system over time.
@@ 4.2 Comparison of Hybrid-AuON and Muon on FineWebEDU-10B
We conduct a controlled paired comparison between Temperature-scaled Hybrid-AuON and Muon on a H100 GPU with a 12-layer GPT-style decoder with hidden size $d=768$ and 6 attention heads(168 Million parameters). Training data are streamed from the 10B-token FineWebEDU kydlicek2025finepdfs corpus using a custom generator that samples contiguous 8k-token sequences while respecting document boundaries via alignment to the GPT-2 end-of-sequence token (50256). Each training step consumes a single 8k-token sequence for both input and target. Validation uses the same sequence length with a fixed budget of 8 sequences (64k tokens), evaluated every 100 steps.
@@
yielding super-linear suppression that engages earlier under spiky gradients.It just change how hard and when the brake engages